ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 184, 2011

Abrams, G., K. M. Rangaswamy and M. Siles Molina The socle series of a Leavitt path algebra

413

Albiac, F. and C. Leránoz Uniqueness of unconditional bases in nonlocally convex c_{0}-products

79

Arveson, W. The noncommutative Choquet boundary II: hyperrigidity

349

Astashkin, S. V. and F. A. Sukochev Symmetric quasi-norms of sums of independent random variables in symmetric function spaces with the Kruglov property

455

Bernig, A. Integral geometry under G_{2} and Spin(7)  301

Billera, L. J. and F. Brenti Quasisymmetric functions and Kazhdan–Lusztig polynomials

317

Brenti, F. See L. J. Billera Calderón, A., A. Kaidi, C. Martín, A. Morales, M. Ramírez

and A. Rochdi Finite-dimensional absolute-valued algebras

193

Cheng, L. Erratum to: Ball-covering property of Banach spaces  505

Dranishnikov, A. N., M. G. Katz and Y. B. Rudyak Cohomological dimension, self-linking, and systolic geometry

437

Elkin, M. An improved construction of progression-free sets  93

Ferreira, R. Quenching phenomena for a non-local diffusion equation with a singular absorption

387

Fish, A. Solvability of linear equations within weak mixing sets  477

Flores, R. J. and R. M. Foote The cellular structure of the classifying spaces of finite groups

129

Foote, R. M. See R. J. Flores

Gismatullin, J. Model theoretic connected components of groups  251

Kaidi, A. See A. Calderón Katz, M. G. See A. N. Dranishnikov Leránoz, C. See F. Albiac Liu, Huifang See Daochun Sun Martín, C. See A. Calderón Martín, M., J. Merí and M. Popov On the numerical index of real

L_{p}(μ)-spaces 183

Merí, J. See M. Martín Mo, X. On some projectively flat Finsler metrics in terms of hyperge-

ometric functions

59

Morales, A. See A. Calderón Popov, M. See M. Martín Rainer, A. Smooth roots of hyperbolic polynomials with definable

coefficients

157

Ramírez, M. See A. Calderón Rangaswamy, K. M. See G. Abrams Rochdi, A. See A. Calderón Rudyak, Y. B. See A. N. Dranishnikov

Ruette, S. Rotation set for maps of degree 1 on the graph sigma  275

Siles Molina, M. See G. Abrams Sukochev, F. A. See S. V. Astashkin Sun, Daochun and Huifang Liu Normal criteria on the family of

meromorphic functions with shared set

403

Usher, M. Boundary depth in Floer theory and its applications to Hamiltonian dynamics and coisotropic submanifolds

1

Vrećica, S. T. and R. T. Živaljević Fulton–MacPherson compactification, cyclohedra, and the polygonal pegs problem

221

Živaljević, R. T. See S. T. Vrećica